Shift report 20030317 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Mon Mar 17 2003 - 07:45:31 EST

  • Next message: Apache: "Shift report 20030317 08:00-16:00"
    ------------------------------------------------------------
    Date: Mon Mar 17 07:45:31 2003, Shift: 00:00-08:00
    Supervisor: Michael Murray, with: Svein Lineal
    Log book#14 pages 66 - 66
    
    Shift Summary:
    ------------------------------------------------------------
    1) We had 2hours and twenty minutes of good store at the beggining of the shift, since then nothing. Currently we are
    in the "machine setup" phase.
    2. We took 3 data runs as well as pedestal and DVM runs.
    3. No problems except that the air conditioner for the FEH hut is broken. Therefore we have turned off the NIM 
    electronics. With the NIM on and an outside temperature of
    49 degrees farenheit the inside temperature was 72 near the door. The high for today is supposed to be 61. Therefore we
    should expect at least 84 degrees in the FEH hut. 
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Mon Mar 17 2003 - 07:46:23 EST